OpenStudy (dilparqq):

In this scale, the frequency of each C-note in vibrations per second is measured in Hz. It can be expressed as a power of 2, as shown. What is the frequency of the note C7 in exponential form and in logarithmic form?

OpenStudy (unklerhaukus):

Middle C = C4 = 261.6 Hz

OpenStudy (unklerhaukus):

C5 = 2 * C4 = 523.2 Hz

OpenStudy (unklerhaukus):

C7 = C4 * 2^3 =